OpenStudy (anonymous):

at 3:00pm catanya and chad had alraedy made 46 cnadles.By 5:00pm, the total reached 100 candles. Assuming a constant production rate,at what time did they make their 82nd candle?

OpenStudy (anonymous):

That's a funny answer but it's 4:20

OpenStudy (anonymous):

how do I work this problem in case I have another question like this?

OpenStudy (anonymous):

YOu have to subtract the hundred and the 46 and then divide that by 120 min(thats the time between 3 and5) so you get a rate of .45 candle/min. Then you subtract the 80 and the 46 and you divide that number by the rate which is .45. Then yo will have 80 and you add that 80min to the 3:00 time to get 4:20

OpenStudy (anonymous):

so 100-46=54 then time between 3:00-5:00 is 120 min. 54/120=.45 82-46=36 36/.45=80min 3:00 + 80min=4:20

OpenStudy (shane_b):

The question can be simplified to the following: Catanya and Chad made 54 candles in 2 hours How long will it take her to make 36 candles? \[\frac{54}{2} = \frac{36}{x}\]Solve for x to see that they made their 82nd candle in 1.333 hours....which is 4:20 pm.